Calculate the final temperature of monatomic ideal gas that is compressed reversibly and adiabatically from 16 litre to 2 litre at 300 Kelvin.Option A 600 KelvinB 1044. 6 KelvinC 1200 KelvinD 750 Kelvin

Grade:12

1 Answers

Tanuj
6 years ago
Let the final Temperature be T. (gamma=5/3)Now we can apply TV^(gamma-1)=constant.300(16)^2/3=T(2)^2/3(8)^2/3=T/3004=T/300We have, T =1200K